Circular No. 7805

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002an IN NGC 2575 S. Nakano, Sumoto, Japan, reports the discovery by Yasuo Sano, Nay-

oro, Hokkaido, of an apparent supernova (V = 16.1) on six CCD frames taken on Jan. 22.52 UT with a 0.28-m f/6.3 Schmidt-Cassegrain reflector. The new object is located at α = 8^{h}22^{m}47^{s}.76, δ = +24^{o}17′41′′.7 (equinox 2000.0), which is 38′′.3 east and 7′′.4 south of the center of NGC 2575. Nothing was visible at this position on CCD frames taken on 1999 Feb. 16 and 2001 Jan. 4 (limiting mag 17.5). R. Kushida, Yatsugatake South Base Observatory, confirmed the presence of the new star at mag 16.0 on an unfiltered CCD image taken on 2002 Jan. 22.64 UT with a 0.40-m f/5.0 Schmidt-Cassegrain reflector, with Y. Kushida providing position end figures 47^{s}.75, 41′′.4. Nothing is present at this location on a survey image taken by R. Kushida on 2002 Jan. 6 (limiting mag 18.5). The Kushidas report SN 2002an at V = 15.9 on Jan. 24.729.

SUPERNOVAE Further to IAUC 7801, B. Barris reports the following additional appar-

ent supernovae on Subaru telescope images taken by P. Challis, K. Chambers, E. Martin, and M. Connelley; most of the new objects were identified by subtracting images taken on 2001 Dec. 9 from those taken on Oct. 22, Nov. 18, and 19. Additional photometric observations were made with Subaru on Jan. 9, the Magellan telescope (+ DirectCam) by S. Jha and R. Kirshner on Dec. 4 and 5, the European Southern Observatory 8.2-m Melipal UT3 Very Large Telescope (+ FORS1) by B. Leibundgut and J. Sollerman on Dec. 13, and with the Canada-France-Hawaii Telescope (+ 12K Mosaic in queue mode) on Dec. 18–21. The magnitudes I_{1} are for the discovery date tabulated; I_{2} below is for Dec. 18 (except Dec. 13 for SN 2001iu and 2002 Jan. 9 for SN 2001ja). The objects below were not present in images taken on Nov. 18 and 19 (limiting mag 25.8).

SN  2001 UT  α_{2000}  δ_{2000}  I_{1}  I_{2}  z
2001iu Dec. 9 7^{h}48^{m}13.^{s}97 +10^{o}15′04′′.5 24.7 24.8
2001iv Dec. 9 7 50 13.53 +10 17 10.4 22.3 22.4 0.4
2001iw Dec. 9 7 50 39.32 +10 20 19.1 22.1 22.5 0.34
2001ix Dec. 9 10 52 18.92 +57 07 29.6 23.6 22.9 0.72
2001iy Dec. 9 10 52 24.28 +57 16 36.1 23.0 23.0 0.567
2001iz Dec. 9 10 53 37.59 +57 15 10.7 23.0 23.0
2001ja Dec. 18 10 52 26.71 +57 19 54.7 23.2 24.1
2002 January 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green